OpenStudy (anonymous):

find the domain x / x^2+25

OpenStudy (anonymous):

All Real numbers

jimthompson5910 (jim_thompson5910):

Set the denominator x^2+25 equal to zero and solve x^2+25=0 x^2=-25 x=sqrt(-25) or x=-sqrt(-25) Since we cannot take the square root of a negative number, this means that there are no real solutions to this equation So the domain is the set of all real numbers.

OpenStudy (anonymous):

jim agrees with me

jimthompson5910 (jim_thompson5910):

chess you're thinking of x^2-25 instead

OpenStudy (anonymous):

Eh, sorry, I misread the sign.

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Let me just delete that....

OpenStudy (anonymous):

lol it's okay

OpenStudy (anonymous):

so its infinity?

jimthompson5910 (jim_thompson5910):

Domain in interval notation is \[\large (-\infty,\infty)\]